OpenStudy (anonymous):

-At the doctor's office, there is a scale that measures people's weight. The needle is 20 cm long and rotated 100° when Frank stepped on it. How far did the needle travel as it rotated?

OpenStudy (anonymous):

9 pi over 100 100 pi over 9 2000π 1000π

OpenStudy (anonymous):

Length of an Arc: \[\frac{ n }{ 360 } \times 2\pi r\] Where n is the angle and r is the radius. In this case n would be the degrees that the needle rotates and the radius is the length of the needle.

OpenStudy (anonymous):

so then it would be b @wotseit

OpenStudy (anonymous):

Correct! Good job :)